# Threshold Graphs are Globally Synchronizing

### v2

@@ Abstract.
The Kuramoto model describes phase oscillators on the unit circle whose interactions are encoded by a graph. Each edge acts like a spring that pulls the two adjacent oscillators toward each other whenever their phases differ. A central question is to determine which graphs are globally synchronizing, meaning that trajectories of the Kuramoto dynamics converge to the fully synchronized state from almost all initial conditions, except for a set of measure zero. This property is tightly linked to the benign nonconvexity of the model’s energy landscape. Existing guarantees for global synchronization rely on minimum-degree thresholds, which require the graph to be highly dense. In this work, we show that connected threshold graphs, whose density may vary from $2/n$ to $1$, are globally synchronizing. Our proof relies on a phasor-geometric analysis of the stationary points of the associated energy landscape.
@@ 1. Introduction
The study of synchronization traces back to the 17th century, when Christiaan Huygens first observed the spontaneous synchronization of pendulum clocks. Since then, synchronization phenomena have been widely studied across both natural and technological systems. For example, the flashing of fireflies [Buck1968, Strogatz1993], the rhythmic beating of heart cells [Winfree1967, Mikhailov2002], synchronization in power grids [dorfler2012synchronization], coordinated motion in robotic networks [OlfatiSaber2006], modern artificial intelligence models such as neural networks [KuramotoNeurons2024], and Transformer architecture [Geshkovski2025, Criscitiello2024].
A major obstacle in the study of synchronization was the absence of a model that simultaneously offered analytical tractability and sufficient dynamical richness. The Kuramoto model, introduced by Yoshiki Kuramoto in 1975 [Kuramoto1975], provided precisely such a balance, and has since become a central and widely accepted model for analyzing synchronization phenomena.
@@
Early studies of the Kuramoto model primarily considered systems of oscillators with heterogeneous intrinsic frequencies $\{\omega_{i}\}_{i=1}^{n}$, coupled through a complete graph with uniform weights $\boldsymbol{A}_{ij}=K/n$. Here, $K$ denotes the strength of the coupling between oscillators. A fundamental question in this setting is to determine the critical coupling strength $K_{c}$, such that for $K>K_{c}$, the oscillators converge to a common frequency, a phenomenon known as frequency synchronization [^1] That is, all derivatives $\tfrac{d\theta_{i}(t)}{dt}$ become identical as $t\to\infty$., and for $K<K_{c}$, they remain desynchronized. This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [michaels1987sinoatrial], technological applications such as deep brain stimulation [tass2003desynchronizing, nabi2011singleinput] and power systems [chiang1995direct, dorfler2012synchronization]. For comprehensive surveys of this line of work and its applications, we refer to [strogatz2000, acebron2005, dorfler2014].
@@
$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in[n].$ (2)
By LaSalle’s invariance principle, every trajectory converges to the invariant set of equilibria satisfying $\dot{\theta}_{i}=0$ for all $i$. Thus, in the co-rotating frame, the asymptotic frequency of every oscillator is zero. Transforming back to the original coordinates, this implies that the oscillators always achieve frequency synchronization. In contrast, when intrinsic frequencies are non-identical, the dynamics may exhibit significantly more complex long-term behavior, including chaotic attractors [Maistrenko2005].
@@
Over the past two decades, a central question that has emerged is the following: under what graph structures does the homogeneous Kuramoto system achieve phase synchronization for all initial conditions, except for a set of measure zero? This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [bandeira2016low, ling2019landscape, ling2023local] and the dynamics of Transformer architectures [Geshkovski2025, geshkovski2023emergence].
@@
We say that a graph $G=(V,E)$ is globally synchronizing if, for almost all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$, the solution to (2) satisfies
@@
Historically, Verwoerd and Mason [verwoerd2007global] conjectured in 2007 that every complete graph $K_{n}$ ($n\geq 2$) is globally synchronizing. This conjecture was resolved by Richard Taylor in 2012 [taylor2012there], and global synchronization of complete graphs has since become a standard benchmark result in this line of research. For classical results on the Kuramoto model on complete graphs, we refer to [acebron2005, strogatz2000, PikovskyRosenblum2015, Bick2020].
In the same work [taylor2012there], Taylor also initiated what is now known as the extremal combinatorics version of the problem [bandeira2025randomstrasse101openproblems2024], which sparked substantial interest and subsequent developments. A more detailed discussion of these advances will be given in Prior Work.
@@
Clearly, this function is nonconvex. Moreover, understanding the geometry of (4) provides direct insight into the long-term dynamics of (2). In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minimua) correspond to non-synchronous stable equilibria. If such spurious local minima exist, the gradient dynamics may converge to them whenever the initialization lies within their basin of attraction, in which case the graph $G$ is not globally synchronizing. In fact, if one can show that the energy function (4) admits no spurious second-order stationary point (in such case we say the function is benign), then the graph $G$ is globally synchronizing. A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality [Lageman2007] together with the center manifold theorem; see Lemma A.1 in [Geshkovski2025] for an excellent exposition.
Prior Work. Existing guarantees for global synchronization on deterministic graphs focus on the extremal combinatorics version of the problem. The aim is to identify minimum-degree thresholds above which a graph is guaranteed to be globally synchronizing.
@@
Let $\delta(G)$ denote the minimum degree of $G$. Find a constant $\mu_{c}\in[0,1]$ such that
$\delta(G)\ \geq\ \mu_{c}(n-1)\ \Longrightarrow\ G\ \text{is globally synchronizing},$
whereas for every $\varepsilon>0$ there exist graphs with $\delta(G)\leq(\mu_{c}-\varepsilon)(n-1)$ that are not globally synchronizing.
The intuition behind this problem is that when every node is connected to a sufficiently large fraction of the network, the graph becomes dense enough for the collective attractive forces to guarantee convergence to synchrony. A remarkable property of these degree-based criteria is that, although they are only sufficient, they apply any graph meeting the threshold. This means global synchronization is guaranteed regardless of the graph’s finer structural details.
Several advances have been made. The first result, due to Taylor [taylor2012there] in 2012, established a minimum-degree threshold of 0.9375(n1)0.9375(n-1). Subsequently, Ling [ling2019landscape] proved that global synchronization holds whenever the minimum degree exceeds 0.7929(n1)0.7929(n-1), and Lu [lu2021synchronization] further improved this bound to 0.7889(n1)0.7889(n-1). The conjecture asserts that the minimum-degree threshold for global synchronization is exactly $3/4$, and below this density, even very dense graphs may fail to synchronize. We refer to Conjecture 5 (density threshold for global synchrony) in [bandeira2025randomstrasse101openproblems2024] for a formal statement of this problem.
@@
Global synchrony on random graphs constitutes another important line of inquiry. Ling, Xu, and Bandeira [ling2019landscape] proved that for Erdős–Rényi graphs, global synchronization occurs with high probability as $n\to\infty$ provided that the edge probability satisfies $p\gg\log(n)/n^{1/3}$. This bound was later improved by Kassabov, Strogatz, and Townsend to $p\gg\log^{2}(n)/n$ [kassabov2022global]. Both works raised the conjecture that synchronization should already occur as soon as the graph is connected, that is, for any $\epsilon>0$ and $p\geq(1+\epsilon)\log(n)/n$. This conjecture was recently resolved in [abdalla2024expandergraphsgloballysynchronizing]. Further improvement has been made in [jain2025random, mcrae2025benign]. For results on global synchrony of random geometric graphs, see [abdalla2024guarantees].
Interestingly, recent work has pointed out that the idealized transformer dynamics admits a similar optimization form as the Kuramoto energy function [Geshkovski2025]. This gives rise to the so-called nonlinear setting of the Kuramoto model, obtained by placing a nonlinear function on top of the inner product $E_{f}(\boldsymbol{\theta})=\-\sum_{1\leq i,j\leq n}A_{ij}f\big(\cos(\theta_{i}-\theta_{j})\big).$ The classical Kuramoto case corresponds to the linear function $f(t)=t$, while in the transformer dynamics one takes $f(t)=e^{\beta t}$ where $\beta>0.$ In this setting, the central question is to determine the values of $\beta$ for which global synchronization occurs, that is, the tokens collapse into a single cluster. Equivalently, this means that all token are making uniform predictions. A well-known conjecture predicted that global synchronization holds for all $\beta\geq 0$. The problem was originally formulated in [Geshkovski2025], with an initial result for $\beta\geq\Omega(1/n)$. It was subsequently improved in [Criscitiello2024] to cover the range $\beta\in[0,1]$, and was very recently settled by Polyanskiy, Rigollet, and Yao [polyanskiy2025synchronization]. For more work connecting the Kuramoto model and Transformer (or attention) dynamics, see [geshkovski2023emergence, karagodin2024clustering, chen2025quantitative, polyanskiy2025synchronization, rigollet2025mean, Geshkovski2024] and the references therein.
Our Contribution. Existing sufficient conditions for global synchronization of deterministic graphs require the graph to be sufficiently dense. In this work, we present a class of graphs that does not satisfy this minimum-degree requirement and whose densities range from 0 to $1$, yet globally synchronizes nevertheless. Through this result. Our main result is the following:
@@
In the following, we briefly introduce the definition of threshold graphs and the motivation for studying their global synchronization. We conclude this section with a discussion of the structural mechanism that underlies the global synchronization of this graph class.
(1) What Are Threshold Graphs? Threshold graphs can be constructed starting from a single vertex by iteratively adding each new vertex as either: an isolated vertex connecting to no existing vertex, or a dominating vertex connecting to all existing vertices. This generative process is uniquely encoded by a binary sequence of length n1n-1, where 0 denotes an isolated vertex and 1 a dominating one. Note that the binary sequence of any connected threshold graph must end with a 1; otherwise, the graph is not connected since there exists at least one isolated node. A full formal definition and additional characterizations are given in Section 2.
@@
(2) Why Threshold Graphs? Many theorems begin with initial observations from which suitable conjectures are formulated, and our case is no exception. The investigation into global synchronization on threshold graphs began with several simple observations:
Observation 1. Star graphs and complete graphs are respectively the densest and sparsest connected threshold graphs, and both exhibit global synchronization.
$\bigl(\underbrace{\texttt{0}\;\texttt{0}\;\cdots\;\texttt{0}}_{\text{all 0}},\;\texttt{1}\bigr)$ $\bigl(\underbrace{\texttt{1}\;\texttt{1}\;\cdots\;\texttt{1}}_{\text{all 1}}\bigr)$
Star graph $S_{n}$ Complete graph $K_{n}$
Certainly, this is not sufficient to propose the conjecture that all threshold graphs globally synchronize. Indeed, although the two extreme cases synchronize, this does not by itself imply that the entire spectrum of threshold graphs does. However, two further observations strengthened our confidence in this conjecture.
Observation 2. One of the many equivalent definitions of threshold graphs requires that their neighborhood-inclusion preorder is total, i.e., the neighborhoods of any two nodes are nested: for any pair $u,v$, either $N(u)\subseteq N[v]$ or $N(v)\subseteq N(u)$. This total preorder of vicinities [threshold_graphs_book] suggests that influence may propagate step by step from the most dominant node to the rest without interference, hinting at the possibility of global synchronization.
Observation 3. While far from exhaustive, our numerical experiments did not reveal any non-synchronous stable equilibria on connected threshold graphs. Our experiments are twofold: (a) To examine global synchronization precisely in small systems, we employed computational algebraic geometry following [mehta2015algebraic]. More explicitly, we reformulated the equilibrium equations of the Kuramoto model as a system of polynomial equations and used the Bertini software package [BHSW13] to compute all solutions. We then evaluated the Jacobian at each solution to determine its stability. (b) For larger size systems (up to 200), using ode45 in Matlab, we initialized a large number of random initial conditions and observed that all trajectories converged to the fully synchronized state.
(3) Differences Between Our Approach and Existing Techniques. Many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma. This lemma states that if all oscillators are confined to an open half-circle, then their dynamics will contract toward synchrony. Based on this observation, the bootstrap strategy attempts to show that all second-order stationary points must lie within a half-circle. If successful, this rules out the existence of non-synchronous state that the dynamics may converge to and thereby implies global synchronization.
However, for threshold graphs, it remains unclear how to apply this strategy. It typically depends on sufficient connectivity in the network thus naturally suited to dense graphs, where the large number of edges ensures that nodes inside the half-circle collectively exert sufficient force to pull outside nodes back into alignment. In contrast, threshold graphs span the full spectrum from extremely sparse examples such as stars to dense cliques, and their highly inhomogeneous, layered structure complicates the application of this strategy. As a result, existing techniques offer limited insight into whether and why these graphs synchronize globally.
@@
(4) Understanding the Mechanism Behind Global Synchrony in Threshold Graphs. Previous global synchronization results focus on density-based conditions. For instance, any graph whose minimum degree exceeds 0.7889(n1)0.7889(n-1) is globally synchronizing [lu2020synchronization]. In other words, these results require the graph to be sufficiently dense. Threshold graphs, however, lie outside this regime. Depending on the positions of dominating vertices in the generating sequence, their density can vary from $O(2/n)$ (stars) to $1$ (cliques), and many intermediate densities occur. Nevertheless, we prove in this work that they also exhibit global synchronization. What deeper insight can we draw from this result?
Our analysis shows that the global synchronization of threshold graphs is not a consequence of density. Instead, it arises from a different mechanism, namely the presence of local symmetries. Some of these symmetries come from the graph structure itself: as we show in Lemma 5.2, closed twins must synchronize at any second-order stationary point. There are also symmetries that do not originate purely from combinatorial structure, that is, certain configurations exhibit a form of geometric symmetry. As we will show in Section 5.3, in the configurations we call synchronous pendants, pairs of vertices that are not structurally symmetric nonetheless become geometrically equivalent at second-order stationary points and hence synchronize. Such symmetric configurations arise inductively along the construction process of threshold graphs. Their synchronization propagates backward from later-added vertices to earlier ones, which shows that every second-order stationary point is a synchronous state. Consequently, every threshold graph is globally synchronizing.
Notation. Throughout this work, we consider only connected simple graphs, i.e., graphs without self-loops or weighted edges. Following the convention of dynamical systems, we refer to the variable $\theta=(\theta_{1},\cdots,\theta_{n})$ of the Kuramoto model (2) as a state, and call the solution $\boldsymbol{\theta}(t)$ with respect to the initial condition $\boldsymbol{\theta}(0)$ in the Kuramoto model (2) a trajectory starting from $\boldsymbol{\theta}(0)$. Moreover, $\mathbb{S}^{1}:=\{x\in\mathbb{R}^{2}:\|x\|=1\}$ denotes the unit circle. And $\boldsymbol{x}\Uparrow\boldsymbol{y}$ means that $\boldsymbol{x}$ and $\boldsymbol{y}$ are nonzero and positively collinear: $\boldsymbol{x}=\lambda\boldsymbol{y}$ for some $\lambda>0$. We write $u\sim v$ to indicate that vertices $u$ and $v$ are adjacent in a graph, and $u\nsim v$ otherwise. For a vertex $i$ and a subset $Q$ of vertices, we denote

### v4

@@ Abstract.
The Kuramoto model describes phase oscillators on the unit circle whose interactions are encoded by a graph. Each edge acts like a spring that pulls adjacent oscillators toward each other whenever their phases differ. A central question is to determine which graphs are globally synchronizing, meaning that trajectories of the Kuramoto dynamics converge to the fully synchronized state from almost all initial conditions. This property is tightly linked to the benign nonconvexity of the model’s energy landscape. Existing guarantees for global synchronization typically require every node to connect to sufficiently many neighbors, which in turn enforces rather homogeneous degree sequences and highly dense graphs. In this work, we identify threshold graphs as a class of graphs that lie outside this regime yet nevertheless exhibit global synchronization. More precisely, connected threshold graphs realize all admissible edge densities ranging from $2/n$ to $1$, and their degree sequences are maximally left-skewed among all graphs with the same edge density, when ordered in nonincreasing order. Our proof relies on a phasor-geometric analysis of the stationary points of the associated energy landscape.
@@ 1. Introduction
The study of synchronization traces back to the 17th century, when Christiaan Huygens first observed the spontaneous synchronization of pendulum clocks. Since then, synchronization phenomena have been widely studied across both natural and technological systems. For example, the flashing of fireflies [BB68, SS93], the rhythmic beating of heart cells [Win67, MS06], synchronization in power grids [DB12], coordinated motion in robotic networks [OS06], modern artificial intelligence models such as neural networks [WCP+24], and Transformer architecture [GLPR25, CRMB24].
A major obstacle in the study of synchronization was the absence of a model that simultaneously offered analytical tractability and sufficient dynamical richness. The Kuramoto model, introduced by Yoshiki Kuramoto in 1975 [Kur75], provided precisely such a balance, and has since become a central and widely accepted model for analyzing synchronization phenomena.
@@
Early studies of the Kuramoto model primarily considered systems of oscillators with heterogeneous intrinsic frequencies $\{\omega_{i}\}_{i=1}^{n}$, coupled through a complete graph with uniform weights $\boldsymbol{A}_{ij}=K/n$. Here, $K$ denotes the strength of the coupling between oscillators. A fundamental question in this setting is to determine the critical coupling strength $K_{c}$, such that for $K>K_{c}$, the oscillators converge to a common frequency, a phenomenon known as frequency synchronization [^1] That is, all derivatives $\tfrac{d\theta_{i}(t)}{dt}$ become identical as $t\to\infty$., and for $K<K_{c}$, they remain desynchronized. This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [MMJ87], technological applications such as deep brain stimulation [Tas03, NM11] and power systems [CCC95, DB12]. For comprehensive surveys of this line of work and its applications, we refer to [Str00, ABPV+05, DB14].
@@
$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V.$ (2)
By LaSalle’s invariance principle, every trajectory converges to the invariant set of equilibria satisfying $\dot{\theta}_{i}=0$ for all $i$. Thus, in the co-rotating frame, the asymptotic frequency of every oscillator is zero. Transforming back to the original coordinates, this implies that the oscillators always achieve frequency synchronization. In contrast, when intrinsic frequencies are non-identical, the dynamics may exhibit significantly more complex long-term behavior, including chaotic attractors [MPT05].
@@
Over the past two decades, a central question that has emerged is the following: under what graph structures does the homogeneous Kuramoto system achieve phase synchronization for all initial conditions, except for a set of measure zero? This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [BBV16, LXB19, Lin23] and the dynamics of Transformer architectures [GLPR25, GLPR23].
@@ 1.1. Problem Setting
We say that a graph $G=(V,E)$ is globally synchronizing if, for all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$ except for a set of measure zero, the solution to the Kuramoto model (2) on $G$ satisfies
@@
Historically, Verwoerd and Mason [VM07] conjectured in 2007 that every complete graph $K_{n}$ ($n\geq 2$) is globally synchronizing. This conjecture was resolved by Richard Taylor in 2012 [Tay12], and global synchronization of complete graphs has since become a standard benchmark result in this line of research. For classical results on the Kuramoto model on complete graphs, we refer to [ABPV+05, Str00, PR15, BGLM20].
In the same work [Tay12], Taylor also initiated what is now known as the extremal combinatorics version of the problem [BKMR25], which sparked substantial interest and subsequent developments. A more detailed discussion of these advances will be given in Prior Work.
@@
Clearly, this function is nonconvex. Moreover, understanding the geometry of (4) provides direct insight into the long-term dynamics of (2). In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minimua) correspond to non-synchronous stable equilibria. If such spurious local minima exist, the gradient dynamics may converge to them whenever the initialization lies within their basin of attraction, in which case the graph $G$ is not globally synchronizing. In fact, if one can show that the energy function (4) admits no spurious second-order stationary point (in such case we say the function is benign), then the graph $G$ is globally synchronizing. A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality [Lag07] together with the center manifold theorem; see Lemma A.1 in [GLPR25] for an excellent exposition.
1.2. Prior Work
Previous works on globally synchronizing graphs primarily focus on identifying sufficient conditions under which a graph is guaranteed to exhibit global synchronization. A prominent line of research formulates this question in an extremal combinatorics version:
@@ 1.2. Prior Work
Let $\delta(G)$ denote the minimum degree of $G$. Find the smallest constant $\mu_{c}\in[0,1]$ such that as long as $\delta(G)\ \geq\ \mu_{c}(n-1)$, the graph is globally synchronizing.
The intuition for this problem is as follows. It is a classical result that complete graphs admit global synchronization [Tay12]. In this case, each node is connected to all other n1n-1 nodes, corresponding to the maximal value $\mu_{c}=1$. A natural question is whether the global synchronization property persists when $\mu_{c}=1$ is relaxed, that is, when each node is connected to only a fraction of the network. Several advances have been made in [Tay12, LXB19, LS21, KST21, Can22, YTT21]. The currently best known bounds on the critical minimum-degree threshold $\mu_{c}$ satisfy $0.6875\leq\mu_{c}\leq 0.75$, where the upper bound is established in [KST21], and the strongest known lower bound follows from [Can22]. It has been conjectured that the minimum-degree threshold for global synchronization is exactly $\mu_{c}=0.75$, and that below this density, even highly connected graphs may fail to synchronize (see [TSS20] and Conjecture 5 in [BKMR25]).
A remarkable feature of these minimum degree-based criteria is that, although they are only sufficient, they apply uniformly to all graphs meeting the prescribed threshold. In particular, global synchronization is guaranteed based solely on this parameter, without invoking any finer structural properties of the graph. Another important observation is that edge density alone is not an appropriate criterion for global synchronization. Indeed, Townsend, Stillman, and Strogatz [TSS20] provided examples of dense graphs that nevertheless fail to synchronize globally. Nevertheless, imposing a minimum degree condition necessarily forces the graph to be very dense. In particular, when $\mu_{c}\geq 75\%$, the total number of edges satisfies
$|E|\geq\frac{3}{8}\,n(n-1),$
which scales as $O(n^{2})$ and is therefore comparable to that of a complete graph.
At the opposite end of the connectivity spectrum from dense networks, how sparse a globally synchronizing graph can be?
Several existing works provide evidence toward this question. First, Townsend, Stillman, and Strogatz [TSS20] approached this problem by asking how few edges are required to transform a ring graph [^3] A ring graph of size $n\geq 5$ does not globally synchronize. into a globally synchronizing network. It was shown that destabilizing all twisted states, which form a family of non-synchronous equilibria of the ring graph on $n$ nodes, requires adding $O(n\log n)$ edges arranged in a specific manner. Numerical experiments further suggest that, for networks of modest size, no non-synchronous stable states persist after such modifications, indicating that certain sparse circulant graphs may exhibit global synchronization.
@@
Second, results on globally synchronizing Erdős–Rényi graphs further indicate that networks substantially sparser than those satisfying the classical $75\%$ minimum-degree conditions can still exhibit global synchronization. Ling, Xu, and Bandeira [LXB19] first proved that global synchronization occurs with high probability as $n\to\infty$ provided that the edge probability satisfies $p\gg\log(n)/n^{1/3}$. This bound was later improved by Kassabov, Strogatz, and Townsend to $p\gg\log^{2}(n)/n$ [KST22]. Both works raised the conjecture that global synchronization should already occur as soon as the graph is connected, namely, for any $\varepsilon>0$ and $p\geq(1+\varepsilon)\frac{\log n}{n}$. This conjecture was recently resolved in [ABK+24], with further refinements obtained in [JMS25, McR25].
Taken together, the sparsest known families of globally synchronizing graphs beyond the trivial case of trees were supported by evidence at the $O(n\log n)$ edge scale. Whether global synchronization can be guaranteed for graphs with significantly fewer edges444beyond the trivial cases tree graphs remained mysterious.
1.3. Our Contribution
In this work, we identify a class of graphs that admit global synchronization without requiring the minimum degree to exceed prescribed thresholds.
@@ 1.3. Our Contribution
Threshold graphs can be constructed starting from a single vertex by iteratively adding each new vertex as either: an isolated vertex connecting to no existing vertex, or a dominating vertex connecting to all existing vertices. This generative process is uniquely encoded by a binary sequence of length n1n-1, where 0 denotes an isolated vertex and 1 a dominating one. See Figure 2 for examples. A full formal definition and additional characterizations are given in Section 2.
@@
Notably, graphs in this class can have arbitrarily low or high edge density, ranging from $2/n$ to $1$, demonstrating that density alone is not an effective indicator of global synchronization. Moreover, their degree sequences are typically highly left-skewed, in sharp contrast to the relatively flat degree sequences arising in prior density-based and random graph models. This comparison further indicates that the shape of the degree sequence, by itself, is also not an appropriate indicator of global synchronization. We discuss the density and degree sequence of threshold graphs in Section 7.
In terms of proving the main result, the difficulty arises from the fact that many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma. This lemma states that if all oscillators are confined to an open half-circle, then the dynamics contract toward synchrony. Building on this observation, the bootstrap strategy seeks to show that all second-order stationary points must lie within a half-circle. If successful, this excludes the existence of non-synchronous states to which the dynamics may converge, thereby implying global synchronization. However, for threshold graphs, it remains unclear how to apply this strategy. It typically depends on sufficient connectivity in the network thus naturally suited to dense graphs, where the large number of edges ensures that nodes inside the half-circle collectively exert sufficient force to pull outside nodes back into alignment. In contrast, threshold graphs span the full spectrum from extremely sparse examples such as stars to dense cliques, and their highly inhomogeneous, layered structure complicates the application of this strategy. As a result, existing techniques offer limited insight into whether and why these graphs synchronize globally.
@@
Understanding the Mechanism Behind Global Synchrony in Threshold Graphs. Previous global synchronization results focus on density-based conditions. For instance, any graph whose minimum degree exceeds 0.7889(n1)0.7889(n-1) is globally synchronizing [LS20]. In other words, these results require the graph to be sufficiently dense. Threshold graphs, however, lie outside this regime. Depending on the positions of dominating vertices in the generating sequence, their density can vary from $2/n$ (stars) to $1$ (cliques), and many intermediate densities occur. Nevertheless, we prove in this work that they also exhibit global synchronization.
What deeper insight can we draw from this result? Our analysis shows that the global synchronization of threshold graphs is not a consequence of density. Instead, it arises from a different mechanism, namely the presence of local symmetries. Some of these symmetries come from the graph structure itself: as we show in Lemma 5.2, closed twins must synchronize at any second-order stationary point. There are also symmetries that do not originate purely from combinatorial structure, that is, certain configurations exhibit a form of geometric symmetry. As we will show in Section 5.3, in the configurations we call synchronous pendants, pairs of vertices that are not structurally symmetric nonetheless become geometrically equivalent at second-order stationary points and hence synchronize. Such symmetric configurations arise inductively along the construction process of threshold graphs. Their synchronization propagates backward from later-added vertices to earlier ones, which shows that every second-order stationary point is a synchronous state. Consequently, every threshold graph is globally synchronizing.
1.4. Structure of the paper
The preliminaries on threshold graphs and the Kuramoto model are presented in Sections 2 and 3, respectively. The basic geometric tools, referred to as phasor geometry or vector geometry, are introduced in Section 4. The key lemmas, termed Local Synchronization Primitives, which underpin our main proof, are presented in Section 5. The proof of the main result is given in Section 6. In Section 7, we discuss the density and degree sequence of threshold graphs. Finally, Section 8 concludes the paper with a discussion of open problems.
1.5. Notation
Throughout this work, we consider only connected simple graphs, i.e., graphs without self-loops or weighted edges. Following the convention of dynamical systems, we refer to the variable $\theta=(\theta_{1},\cdots,\theta_{n})$ of the Kuramoto model (2) as a state, and call the solution $\boldsymbol{\theta}(t)$ with respect to the initial condition $\boldsymbol{\theta}(0)$ in the Kuramoto model (2) a trajectory starting from $\boldsymbol{\theta}(0)$. Moreover, $\mathbb{S}^{1}:=\{x\in\mathbb{R}^{2}:\|x\|=1\}$ denotes the unit circle. And $\boldsymbol{x}\Uparrow\boldsymbol{y}$ means that $\boldsymbol{x}$ and $\boldsymbol{y}$ are nonzero and positively collinear: $\boldsymbol{x}=\lambda\boldsymbol{y}$ for some $\lambda>0$. We write $u\sim v$ to indicate that vertices $u$ and $v$ are adjacent in a graph, and $u\nsim v$ otherwise. For a vertex $i$ and a subset $Q$ of vertices, we denote